# Exploring the dependence of the Hubble constant from the cluster-lensed supernova SN Refsdal on mass model assumptions

### v1

@@ I Introduction
The Hubble constant, denoted as $H_{0}$, quantifies the current expansion rate of the Universe and stands as a key parameter in modern cosmology. The Hubble constant is important, particularly because it is a cornerstone in determining fundamental properties such as the age, size, and fate of the University Hubble (1929); Lemaître (1931); Freedman et al. (2001); Nussbaumer and Bieri (2011); Weinberg et al. (2013). Various methodologies have been employed to measure $H_{0}$. Early estimates relied on distance measurements to relatively nearby galaxies, while contemporary approaches incorporate diverse cosmological probes. These approaches include Type Ia supernova (SNe Ia), cosmic microwave background radiation (CMBR), and baryon acoustic oscillations (BAO), among others. However, despite significant observational progress, reconciling different measurements and reducing systematic uncertainties remain ongoing challenges. The quest to precisely determine $H_{0}$ has led to intriguing controversies Verde et al. (2019); Freedman (2021); Abdalla et al. (2022); Hu and Wang (2023); Linder (2023); Cervantes-Cota et al. (2023), which emerge when comparing local $H_{0}$ measurements with those inferred from observations of the early Universe. For instance, the distance ladder analysis from SH0ES (SNe, H0, for the Equation of State of dark energy) collaboration measures the Hubble constant of $H_{0}=73.04\pm 1.04\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$ in the late Universe Riess et al. (2022), while CMBR from the Planck satellite releases a much smaller and tighter measurement with the value of $H_{0}=67.4\pm 0.5\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$ in the early Universe Planck Collaboration et al. (2020).
@@
Very recently, the Hubble constant has been derived from the cluster-lensed supernova system SN Refsdal in the field of MACS J1149+2223 which is the first reported case of strongly lensed supernovae with resolved multiple images Kelly et al. (2015, 2016). Focusing on this unique and valuable system, Ref. Vega-Ferrero et al. (2018) has carried out Refsdal’s original proposal to use a multiply imaged supernova to measure $H_{0}$, with the value of $H_{0}=64^{+9}_{-11}\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$, in which preliminary values of time delay measurements from the early data in light curve observations have been used. Ref. Kelly et al. (2023a) measure $H_{0}=66.6^{+4.1}_{-3.3}\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$ based on accurate and precise measurements of time delays from the full light curve data. Since their analysis is blinded in the sense that mass models used for the analysis are those constructed before the reappearance of the fifth image of SN Refsdal, they do not thoroughly analyze degeneracies between the lens models and the Hubble constant. Furthermore, independent measurements of the Hubble constant and the geometry of the Universe from SN Refsdal is presented by Ref. Grillo et al. (2024), which argues that the mass model uncertainty is subdominant Grillo et al. (2020).
@@
SN Refsdal as the first example of a strongly lensed supernova with resolved multiple images was initially detected in November 2014 through the Hubble Space Telescope (HST) observations Kelly et al. (2015). This system is composed of six multiple images (named S1-S4, SX, and SY in Oguri (2015)) of a single supernova at redshift 1.488, where four images S1–S4 forming an Einstein-cross configuration around an early-type galaxy were detected by the Grism Lens Amplified Survey from Space (GLASS; GO-13459; PI Treu) Schmidt et al. (2014); Treu et al. (2015), SX detected in HST imaging on 2015 December 11 (GO-14199; PI Kelly) Kelly et al. (2016) was the first supernova whose appearance on the sky was predicted in advance Oguri (2015); Sharon and Johnson (2015); Diego et al. (2016), and a leading image SY appeared in the late 1990s from the lens models predication without observation evidence. SN Refsdal is lensed by the cluster of galaxies MACS J1149.5+2223 at redshift 0.541 Ebeling et al. (2007). Multiple images and their spectroscopic redshifts for this cluster have been confirmed and presented in the literature Zitrin and Broadhurst (2009); Smith et al. (2009); Zheng et al. (2012); Rau et al. (2014); Richard et al. (2014); Jauzac et al. (2016); Treu et al. (2016); Grillo et al. (2016); Brammer et al. (2016).
In this study, we utilize the positions of the five supernova images in Ref. Kelly et al. (2023b), which have been refined based on the Hubble Space Telescope (HST) Wide Field Camera 3 (WFC3) infrared (IR) detector F125W images acquired in 2015 and 2016. Additionally, we incorporate positional constraints derived from multiple images of seven knots within a lensed face-on spiral galaxy at redshift 1.488. For multiple images of other background galaxies, we include data from 28 multiple image systems compiled in Ref. Kawamata et al. (2016). The total number of multiple images in our study is 109, spanning across 36 systems. Our analysis includes newly measured time delays between lensed supernova images S2–SX and S1, as well as recently determined magnification ratios from the full light curve analysis Kelly et al. (2023b). The magnification ratio S4/S1 is excluded due to compelling evidence ($>5\sigma$) indicating that image S4 has undergone chromatic microlensing Kelly et al. (2023a).
@@
Name $\Delta x$ [${}^{\prime\prime}$] $\Delta y$[${}^{\prime\prime}$] $\Delta\mu$ $\Delta t$ [days]
@@ III.1 Dark matter halo
We employ the Navarro-Frenk-White (NFW) profile to characterize the mass distribution of a dark matter halo Navarro et al. (1997). The projected density profile is defined by six parameters, including the mass $M$, the centroid position $x$ and $y$, the ellipticity $e$, the position angle $\theta_{\rm e}$, and the concentration parameter $c$. In our mass model reconstruction, we include up to four NFW components to model the complex cluster mass distribution. For NFW components located at the edge or outside the strong lensing regions, positions are fixed to observed positions of bright cluster member galaxies. In cases where NFW components lack strong observational constraints, the concentration parameter is fixed to 10. Across all halo components, the halo mass and ellipticity are treated as free parameters. We optimize these model parameters, with the ellipticity restricted to be smaller than 0.8 and the concentration parameter restricted to the range between 1 and 40. We utilize the fast approximation of the lensing calculation of the NFW profile, dubbed as anfw in glafic, as proposed in Ref. Oguri (2021).
@@ IV Constraints on Hubble constant
Label Model * $\sigma_{\rm s}$[${}^{\prime\prime}$] $\sigma_{\rm h}$[${}^{\prime\prime}$] $\sigma_{\rm g}$[${}^{\prime\prime}$] $H_{0}$ [$\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$] $\chi^{2}$/dof
@@
Figure 3: Critical curves for the 23 different lens mass models, computed assuming the source redshift of $z_{\rm s}=6$. Left: Critical curves for 12 different lens mass models with best-fitting values of the Hubble constant larger than $70.0\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$. Right: Critical curves for 11 different lens mass models with best-fitting values of the Hubble constant smaller than $70.0\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$. Black squares represent lensed supernova images.
@@ V.1 Critical curves
The shape and extent of critical curves are highly sensitive to the mass distribution of the lensing objects. Therefore, we begin, by checking critical curves, to examine potential dependence of Hubble constant values with cluster mass distributions. The results are displayed in Fig. 3. We find that there are some notable differences of critical curves, especially in the northern part. These differences arise largely from the consideration of varying numbers of dark matter halo components. The large variation of critical curves in the northern part, however, is seen in both models with large and small values of the Hubble constant. Our analysis reveals no significant correlations between critical curve shapes and and the Hubble constant constraints across all examined lens mass models.
@@ V.2 Radial convergence and tangential shear profiles
Next, we calculate the azimuthal average of the convergence at each radius to derive radial convergence and tangential shear profiles for all the 23 different lens mass models. The radial convergence, denoted by $\kappa$, is directly related to the surface mass density of the lensing object and provides insights into its overall mass distribution. On the other hand, the tangential shear, denoted by $\gamma_{\rm T}$, measures the coherent distortion of background sources around the lensing object. Unlike the radial convergence, the tangential shear describes the anisotropic stretching of images along the tangential direction with respect to the lens center. The tangential shear is sensitive to the gradient of the lens potential and offers valuable information about the internal structure and orientation of the lensing mass distribution.
@@ V.3 Magnification factors
Ref. Oguri and Kawano (2003) proposed that the measurement of magnification factors of lensed supernova images, which is available when the source is Type Ia supernova, breaks the degeneracy between the radial density profile and the Hubble constant and significantly improves the constraint on the Hubble constant, because magnification factors are sensitive to the radial density profile. It is however not immediately clear whether this method is indeed useful for massive clusters for which the radial profile is already constrained by many multiple images and also their mass distributions are highly complex. Therefore, it is of great interest to explicitly check whether there is a correlation between best-fitting Hubble constant values and the magnification factors for five multiple images of SN Refsdal, even though SN Refsdal is a core-collapse supernova and hence their magnification factors have not been measured.
@@ VI Conclusion
Our finding appears to differ from our previous work, which revealed significant lens mass model dependence in the cluster-lensed quasar system SDSS J1004+4112 and yielded a weak constraint on the Hubble constant of $H_{0}=67.5^{+14.5}_{-8.9}\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$ Liu et al. (2023). We argue that the large number of multiple images for SN Refsdal plays a crucial role in breaking the degeneracy among different lens mass models and in obtaining the robust measurement of the Hubble constant.
@@ Acknowledgements.
YL was supported by the China Scholarship Council (Grant No. 202106040084). MO was supported by JSPS KAKENHI Grant Numbers JP22H01260, JP20H05856, and JP22K21349.

### v2

# Assessing the effect of mass-model assumptions on measuring the Hubble constant from the cluster-lensed supernova Refsdal

@@ I Introduction
The Hubble constant, denoted as $H_{0}$, quantifies the current expansion rate of the Universe and stands as a key parameter in modern cosmology. The Hubble constant is important, particularly because it is a cornerstone in determining fundamental properties such as the age, size, and fate of the Universe Hubble (1929); Lemaître (1931); Freedman et al. (2001); Nussbaumer and Bieri (2011); Weinberg et al. (2013). Various methodologies have been employed to measure $H_{0}$. Early estimates relied on distance measurements to relatively nearby galaxies, while contemporary approaches incorporate diverse cosmological probes. These approaches include Type Ia supernova (SNe Ia), cosmic microwave background radiation (CMBR), and baryon acoustic oscillations (BAO), among others. However, despite significant observational progress, reconciling different measurements and reducing systematic uncertainties remain ongoing challenges. The quest to precisely determine $H_{0}$ has led to intriguing controversies Verde et al. (2019); Freedman (2021); Abdalla et al. (2022); Hu and Wang (2023); Linder (2023); Cervantes-Cota et al. (2023), which emerge when comparing local $H_{0}$ measurements with those inferred from observations of the early Universe. For instance, the distance ladder analysis from SH0ES (SNe, H0, for the Equation of State of dark energy) collaboration measures the Hubble constant of $H_{0}=73.04\pm 1.04\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$ in the late Universe Riess et al. (2022), while CMBR from the Planck satellite releases a much smaller and tighter measurement with the value of $H_{0}=67.4\pm 0.5\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$ in the early Universe Planck Collaboration et al. (2020).
@@
Very recently, the Hubble constant has been derived from the cluster-lensed supernova system SN Refsdal in the field of MACS J1149.5+2223 which is the first reported case of strongly lensed supernovae with resolved multiple images Kelly et al. (2015, 2016). Focusing on this unique and valuable system, Ref. Vega-Ferrero et al. (2018) has carried out Refsdal’s original proposal to use a multiply imaged supernova to measure $H_{0}$, with the value of $H_{0}=64^{+9}_{-11}\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$, in which preliminary values of time delay measurements from the early data in light curve observations have been used. Ref. Kelly et al. (2023a) measure $H_{0}=66.6^{+4.1}_{-3.3}\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$ based on accurate and precise measurements of time delays from the full light curve data. Since their analysis is blinded in the sense that mass models used for the analysis are those constructed before the reappearance of the fifth image of SN Refsdal, they do not thoroughly analyze degeneracies between the lens models and the Hubble constant. Furthermore, independent measurements of the Hubble constant and the geometry of the Universe from SN Refsdal is presented by Ref. Grillo et al. (2024), which argues that the mass model uncertainty is subdominant Grillo et al. (2020).
@@
SN Refsdal as the first example of a strongly lensed supernova with resolved multiple images was initially detected in November 2014 through the Hubble Space Telescope (HST) observations Kelly et al. (2015). This system is composed of six multiple images (named S1-S4, SX, and SY in Oguri (2015)) of a single core-collapse supernova at redshift 1.488, where four images S1–S4 forming an Einstein-cross configuration around an early-type galaxy were detected by the Grism Lens Amplified Survey from Space (GLASS; GO-13459; PI Treu) Schmidt et al. (2014); Treu et al. (2015), SX detected in HST imaging on 2015 December 11 (GO-14199; PI Kelly) Kelly et al. (2016) was the first supernova whose appearance on the sky was predicted in advance Oguri (2015); Sharon and Johnson (2015); Diego et al. (2016), and a leading image SY appeared in the late 1990s from the lens models predication without observation evidence. SN Refsdal is lensed by the cluster of galaxies MACS J1149.5+2223 at redshift 0.541 Ebeling et al. (2007). Multiple images and their spectroscopic redshifts for this cluster have been confirmed and presented in the literature Zitrin and Broadhurst (2009); Smith et al. (2009); Zheng et al. (2012); Rau et al. (2014); Richard et al. (2014); Jauzac et al. (2016); Treu et al. (2016); Grillo et al. (2016); Brammer et al. (2016).
In this study, we utilize the positions of the five supernova images in Ref. Kelly et al. (2023b), which have been refined based on the Hubble Space Telescope (HST) Wide Field Camera 3 (WFC3) infrared (IR) detector F125W images acquired in 2015 and 2016. Additionally, we incorporate positional constraints derived from multiple images of seven knots within a lensed face-on spiral galaxy at redshift 1.488. For multiple images of other background galaxies, we include data from 28 multiple image systems compiled in Ref. Kawamata et al. (2016). The total number of multiple images in our study is 109, spanning across 36 systems. Fig. 1 shows the position of all the multiple images used for our analysis. Our analysis includes newly measured time delays between lensed supernova images S2–SX and S1, as well as recently determined magnification ratios from the full light curve analysis Kelly et al. (2023b). The magnification ratio S4/S1 is excluded due to compelling evidence ($>5\sigma$) indicating that image S4 has undergone chromatic microlensing Kelly et al. (2023a).
@@
Figure 1: The HST F606W image of MACS J1149.5+2223 with positions of multiple images of SN Refsdal (cyan) as well as multiple images of other background galaxies (magenta). North is up and West is right. Positions of halo components are also indicated by large circles. For red circles, the positions are left as free parameters, while for blue circles the positions are fixed in mass modeling. When including only 3 halos in mass modeling, ’halo 1’ is not included.
@@
Name $\Delta x$ [′′] $\Delta y$[′′] $\Delta\mu$ $\Delta t$ [days]
@@ III.1 Dark matter halo
We employ the Navarro-Frenk-White (NFW) profile to characterize the mass distribution of a dark matter halo Navarro et al. (1997). The projected density profile is defined by six parameters, including the mass $M$, the centroid position $x$ and $y$, the ellipticity $e$, the position angle $\theta_{\rm e}$, and the concentration parameter $c$. In our mass model reconstruction, we include up to four NFW components to model the complex cluster mass distribution. For NFW components located at the edge or outside the strong lensing regions, positions are fixed to observed positions of bright cluster member galaxies (see Fig. 1). In cases where NFW components are located outside the main strong lensing region and lack strong observational constraints, the concentration parameter is fixed to 10, because the concentration parameter value is not very well determined from the data. Across all halo components, the halo mass and ellipticity are treated as free parameters. We optimize these model parameters, with the ellipticity restricted to be smaller than 0.8 and the concentration parameter restricted to the range between 1 and 40. We utilize the fast approximation of the lensing calculation of the NFW profile, dubbed as anfw in glafic, as proposed in Ref. Oguri (2021).
@@ IV Constraints on Hubble constant
Label Model * $\sigma_{\rm s}$[′′] $\sigma_{\rm h}$[′′] $\sigma_{\rm g}$[′′] $H_{0}$ [$\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$] $\chi^{2}$/dof
@@
Figure 4: Critical curves for the 23 different lens mass models, computed assuming the source redshift of $z_{\rm s}=6$. Left: Critical curves for 12 different lens mass models with best-fitting values of the Hubble constant larger than $70.0\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$. Right: Critical curves for 11 different lens mass models with best-fitting values of the Hubble constant smaller than $70.0\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$. Black squares represent lensed supernova images. The lower right of each panel shows more details around the lensed supernova images.
@@ V.1 Critical curves
The shape and extent of critical curves are highly sensitive to the mass distribution of the lensing objects. Therefore, we begin, by checking critical curves, to examine potential dependence of Hubble constant values with cluster mass distributions. The results are displayed in Fig. 4. We find that there are some notable differences of critical curves, especially in the northern part. These differences arise largely from the consideration of varying numbers of dark matter halo components. The large variation of critical curves in the northern part, however, is seen in both models with large and small values of the Hubble constant. This means that the uncertainty of the northern structure of the critical curve on the Hubble constant is not very significant. To summarize, our analysis reveals no significant correlations between critical curve shapes and the Hubble constant constraints across all examined lens mass models.
@@ V.2 Radial convergence and tangential shear profiles
Next, we calculate the azimuthal average of the convergence at each radius to derive radial convergence and tangential shear profiles around the main halo (’halo 2’ in Fig. 1) for all the 23 different lens mass models. The radial convergence, denoted by $\kappa$, is directly related to the surface mass density of the lensing object and provides insights into its overall mass distribution. On the other hand, the tangential shear, denoted by $\gamma_{\rm T}$, measures the coherent distortion of background sources around the lensing object. Unlike the radial convergence, the tangential shear describes the anisotropic stretching of images along the tangential direction with respect to the lens center. The tangential shear is sensitive to the gradient of the lens potential and offers valuable information about the internal structure and orientation of the lensing mass distribution.
@@ V.3 Magnification factors
Ref. Oguri and Kawano (2003) proposed that the measurement of magnification factors of lensed supernova images, which is available when the source is Type Ia supernova, breaks the degeneracy between the radial density profile and the Hubble constant and significantly improves the constraint on the Hubble constant, because magnification factors are sensitive to the radial density profile. It is however not immediately clear whether this method is indeed useful for massive clusters for which the radial profile is already constrained by many multiple images and also their mass distributions are highly complex. Therefore, it is of great interest to explicitly check whether there is a correlation between best-fitting Hubble constant values and the magnification factors for five multiple images of SN Refsdal. We caution that SN Refsdal is a core-collapse supernova and hence their magnification factors have not been measured. As a result, in our measurements of the Hubble constant with SN Refsdal, any information on the magnification factors are not used. Our purpose here is to check the potential usefulness of lensed Type Ia supernovae in future analysis.
@@ VI Conclusion
Our finding appears to differ from our previous work, which revealed significant lens mass model dependence in the cluster-lensed quasar system SDSS J1004+4112 and yielded a weak constraint on the Hubble constant of $H_{0}=67.5^{+14.5}_{-8.9}\mathrm{~{}km~{}s^{-1}~{}Mpc^{-1}}$ Liu et al. (2023). We argue that the much larger number of multiple images for SN Refsdal than for SDSS J1004+4112 plays a crucial role in breaking the degeneracy among different lens mass models and in obtaining the more accurate measurement of the Hubble constant.
@@ Acknowledgements.
We thank anonymous referees for useful comments and suggestions. YL was supported by the China Scholarship Council (Grant No. 202106040084). MO was supported by JSPS KAKENHI Grant Numbers JP25H00662, JP25H00672, and JP22K21349.